A H2SO4 solution has pH=3. its normality will be??

Grade:12th pass

1 Answers

Aarti Gupta
9 years ago
pH = -log[ H]+
So,

· [H+]=10-pH=10-3=.001M

Thus 0.001M H2SO4

Normality = Molarity * Basicity (for acids)

= 0.001 * 2 ( As H2SO4 gives 2H+ ions)

Therefore normality comes out to be 0.002N

Thanks & Regards

Aarti Gupta

askiitians Faculty